Fermi National Laboratory
 

Accelerator Update for September 29, 2006 - October 2, 2006

 

Friday September 29

The day shift began with Operations monitoring a Recycler stash (1.2E10) and the Antiproton Source (Pbar) stack (36.5E10), with the Tevatron (TeV) recovering from a quench, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for colliding beam, and with beam to the Linac, Booster, Main Injector (MI), Recycler, NuMI and MiniBooNE.

The Recycler completed a stash of antiprotons at 10:27 AM.

Operations put the TeV through a dry squeeze at 11:07 AM.

Operations established a 14x0 (the first number is for bunches of protons and the second number is for bunches of antiprotons) study store at 12:32 PM.

Operations began preparing the TeV at 3:59 PM for a D0 access.

The Recycler completed a stash at 4:41 PM.

D0 came out of access at 5:56 PM.

Operations put the TeV into shot setup at 7:16 PM.

At 8:26 PM, a MI power supply (I:SEXHPS) tripped off on a ground fault indication. Operations contacted an expert. This problem put shot setup on hold.

At 10:23 PM, MI power supply experts reported that the magnet appeared to have a ground fault. Operation began preparing MI for an access.

Operations went ahead and used the beam they had already loaded into the TeV to establish store 4989 at 10:59 PM with an initial average luminosity of 17.39E30.

Saturday September 30

The midnight shift began with Operations monitoring store 4989 (luminosity 13.8E30) and the stash and stack, and with MI off due to a possible SEXH ground fault.

At 2:29 AM, MI experts reported that they found a bad splice in the cable tray for SEXH. They began repairs.

MI power supply experts had SEXH up and running by 4:18 AM.

Operations established beam to MiniBooNE, NuMI, and Pbar by 4:53 AM.

The Recycler completed a stash around 5:50 AM and then again at 9:36 AM.

Operation terminated store 4989 at 10:03 AM and then put the TeV into shot set a few minutes later.

Operations established store 4990 at 12:27 PM with an initial luminosity of 85.12E30.

The Recycler completed a stash at 1:59 PM.

At 2:24 PM, store 4990 quenched at TeV sector F4 due to a possible B1 separator spark.

Operations put the TeV through a dry squeeze at 6:06 PM.

The Recycler completed a stash at 6:53 PM.

Operations put the TeV into shot setup at 7:23 PM.

Operations established store 4994 at 9:29 PM with an initial luminosity of 81.3E30.

Sunday October 1

The midnight shift began with Operations monitoring store 4994 and the stash and stack, with the Recycler stashing, and with MiniBooNE and NuMI taking beam.

At 12:09 AM, during a stash, Operations reported a problem with a Recycler power supply (H130) not making it up to its current and causing the loss of transfers. The stash ended at around 12:13 AM. Operations contacted experts.

Power supply experts had H130 repaired by 1:42 AM.

The Recycler completed a stash around 4 AM.

The Recycler completed a stash at 7:59 AM.

At 8:58 AM, Operations reported an ongoing problem with a MI power supply (I:HVF12) that had been causing numerous NuMI beam permit trips. An expert had been and continued to work on the problem.

The Recycler completed a stash at 11:47 AM.

The Recycler completed a stash at 3:15 PM.

A TeV expert began a Tevatron Electron Lens (TEL) two-hour study at 4:04 PM.

Operations terminated store 4994 at 6:03 PM.

Operations prepared CDF for a collision hall access 6:23 PM.

The Recycler completed a stash at 7:32 PM.

Operations put the TeV into a dry squeeze at 7:59 PM and then into shot setup at 8:06 PM.

At 9:47 PM, Operations reported a vacuum burst in the A1 transfer line. The vacuum took about ten minutes to recover.

Operations established store 4996 at 11 PM with an initial luminosity of 172.0E30.

Monday October 2

The midnight shift began with Operations monitoring store 4996 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:33 AM, Operations reported continuing problems with NuMI trips due to HVF12.

The Recycler completed a stash at 1:25 AM.

The Recycler completed a stash at 4:30 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar and MI studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/02/2006    email Fermilab